# CORE: Comprehensive Post-OCR Parsing and Receipt Understanding Dataset

### v1

@@ 6.1 Object Detection Results
Table 6 shows the object classification accuracy for various detection methods across different backbone architectures. CAM reveals strong and consistent performance, with scores of 43.10 for ResNet50 and 43.74 for VGG16. Conversely, the Hide-and-Seek (HAS) method displays significant variability, excelling with a score of 34.15 on ResNet50 but performing poorly with VGG16, where it achieves accuracy of only 16.36. This indicates the dependency of HAS’s effectiveness on the backbone’s feature extraction capability. Advanced methods such as the Attention-based Dropout Layer (ADL) and Adversarial Complementary Learning (ACOL) both perform well, particularly with InceptionV2, each with 42.09, suggesting that methods which diversify the model’s attention across the image can significantly improve classification outcomes in complex datasets like CORU.
Table 3: Performance Metrics for Few-shot Information Extraction Across Multiple Categories
@@
6.2 OCR Results
The OCR task of the CORU dataset is important for information extraction. Table 6 summarizes the Character Error Rate (CER) and Word Error Rate (WER) metrics of various OCR models. As a baseline, Tesseract OCR shows a Character Error Rate (CER) of 15.56% and a Word Error Rate (WER) of 30.78%. While Tesseract is widely used due to its open-source nature and extensive language support, the Attention-Gated CNN-BiGRU model improves upon the baseline, reducing the CER to 14.85% and the WER to 27.22%. This model combines gated convolutional neural networks with bidirectional GRU layers, enhanced by attention mechanisms, which likely contributes to its improved handling of spatial dependencies and contextual information within the text. Our OCR Model, designed specifically for the challenges presented by the CORU dataset, further reduces the CER to 7.83%. However, its WER is comparable to that of the Attention-Gated CNN-BGRU at 27.24%, suggesting that while it is more adept at recognizing individual characters, it struggles similarly at the word level. Azura OCR achieves the best performance among the tested models, with the lowest CER of 6.39% and WER of 25.97%.
Table 4: Performance Comparison of Object Classification Methods Across Different Backbone Architectures.
@@ 6.2 OCR Results
Table 5: OCR Performance Comparison
Model CER (%) WER (%)
Tesseract OCR 15.56 30.78
Attention-Gated 14.85 27.22
Our OCR Model 7.83 27.24
Azura OCR 6.39 25.97
Table 6: Performance Evaluation of YOLO Models
Model P R mAP50 mAP50-95
YoloV7 76.00 85.60 79.20 43.70
YoloV8 74.60 81.00 76.10 45.30
YoloV9 75.70 83.40 77.90 46.70
6.3 Information Extraction Results
The exploration of information extraction capabilities of language models across zero-shot, one-shot, and few-shot scenarios reveals significant insights into their adaptability and effectiveness. In zero-shot learning, models generally struggle with precision and recall in complex categories like ’Brand’ and ’Weight’, with Zephyr and Falcon showing relatively better generalization capabilities. The one-shot scenario indicates improvements across the board, with LLaMA V1 demonstrating considerable robustness. Few-shot learning results, detailed in Table 3, demonstrate enhancements in model performance as they were exposed to more examples, affirming the value of incremental learning in complex extraction tasks. These observations underscore the capabilities of current language models to adapt to the demands of extracting structured information from unstructured datasets like receipts.
7 Conclusion
In this work, we presented CORU, a comprehensive dataset designed for enhancing OCR and post-OCR processing in multilingual contexts, particularly involving Arabic and English texts. Our dataset, comprising over 20,000 annotated receipts, supports a range of computational tasks including object detection, OCR, and detailed information extraction. We introduced several models and evaluated their performance across these tasks, demonstrating also the capabilities of our novel OCR model and various object detection frameworks. The results underscore the potential of CORU to facilitate advanced research and practical applications in document understanding and automation.

### v2

@@ 4. Models
Figure 7. Performance of different models on the Receipt QA subset across key metrics.
Figure 8. OCR Performance Comparison
Table 3. Performance Evaluation of YOLO Models
Model P R mAP50 mAP50-95
YoloV7 76.00 85.60 79.20 43.70
YoloV8 74.60 81.00 76.10 45.30
YoloV9 75.70 83.40 77.90 46.70
5. Experimental Results
5.1. QA Results
Figure 7 provides a comparative analysis of various large language models on the Receipt QA subset of ReceiptSense across four key metrics: precision, recall, exact match, and contains. GPT-4o consistently outperforms other models, achieving the highest scores across all metrics — notably 37.7% precision, 36.4% recall, 35.0% exact match, and 29.1% contains. Llama3.2 (11B) follows, with precision at 32.6%, recall at 31.3%, exact match at 31.6%, and contains at 25.9%. Phi3 (4.15B) and Phi3.5 show solid but slightly lower performance, clustering in the 28–30% range for precision, recall, and exact match. Llava (7.06B) and Internvl2 models (4B, 8B) exhibit notably lower results, especially in the exact match and contains metrics, reflecting challenges in handling complex receipt QA tasks. The figure illustrates the clear advantage of larger and more advanced models like GPT-4o and Llama3.2 in comprehending and extracting accurate information from multilingual receipts.
Table 4. Performance of Language Models in Information Extraction Across Zero-Shot, One-Shot, Two-Shot, and Three-Shot Settings for Various Receipt Information Categories.
@@ 5.1. QA Results
Table 5. Performance Comparison of Object Classification Methods Across Different Backbone Architectures.
@@
5.2. Object Detection Results
We evaluated multiple object detection models including Class Activation Mapping (CAM), Hide-and-Seek (HAS), Attention-based Dropout Layer (ADL), Adversarial Complementary Learning (ACOL), Self-produced Guidance (SPG), CutMix, and DINO across different backbone architectures (ResNet50, VGG16, InceptionV2, Swin Transformer). DINO with Swin 4-scale backbone achieved the best performance with an average score of 32.2 and notable robustness across quality thresholds (45.4 at 10%). Traditional models like CAM and HAS showed limitations, with ResNet50 scoring averages of 6.17 and 7.14 respectively, with significant performance drops at higher quality thresholds.
Object classification results revealed strong performance from CAM with consistent scores (43.10 for ResNet50, 43.74 for VGG16). HAS showed significant variability, excelling with ResNet50 (34.15) but performing poorly with VGG16 (16.36), indicating dependency on backbone feature extraction capabilities. Advanced methods like ADL and ACOL performed well with InceptionV2 (both achieving 42.09), suggesting that attention diversification across images improves classification in complex datasets.
YOLO model evaluation provided insights into real-time detection capabilities. YOLOv7 demonstrated highest performance with 76.00% precision, 85.60% recall, and 79.20% mAP50, showing robustness in detecting objects on multilingual receipts with diverse layouts. YOLOv8 showed slightly lower performance but improved mAP across different IoU thresholds (45.30% mAP50-95), suggesting better generalization across object sizes and shapes. YOLOv9 further improved these metrics with 46.70% mAP50-95, highlighting continuous YOLO architecture advancements.
5.3. OCR Results
As shown in Figure 8, OCR evaluation focused on Character Error Rate (CER) and Word Error Rate (WER) metrics. Tesseract OCR baseline achieved 15.56% CER and 30.78% WER. The Attention-Gated CNN-BiGRU model improved performance with 14.85% CER and 27.22% WER by combining gated CNNs with bidirectional GRU layers and attention mechanisms for better spatial dependencies and contextual information handling. Our specialized OCR model reduced CER to 7.83% but maintained similar WER (27.24%), indicating better individual character recognition while struggling at word level. Azura OCR achieved best performance with 6.39% CER and 25.97% WER.
5.4. Information Extraction Results
In zero-shot settings, models performed poorly, particularly for Brand and Weight categories. Zephyr achieved highest performance (42.98 F1, 27.02% accuracy), while Mistral followed (32.07 F1, 18.53% accuracy). The #Units category showed better results with Mistral obtaining 74.94 F1 and 60.52% accuracy, indicating easier numerical field recognition. One-shot learning showed marked improvement, especially LLaMA V1 in Brand category (jumping from 3.70 to 47.27 F1). Mixtral led Weight extraction (58.07 F1, 40.93% accuracy), while Mistral achieved best overall performance (60.60 F1, 43.56% accuracy). Few-shot learning further enhanced performance: in two-shot scenarios, Zephyr recorded 52.43 F1 and 35.39% accuracy for Brand category, while LLaMA V2 excelled in #Units extraction (93.42 F1, 89.24% accuracy). Three-shot setup showed Falcon leading Brand category (62.30 F1, 45.39% accuracy) and Mistral dominating Weight category (76.11 F1, 62.08% accuracy), with Mistral achieving best overall performance (80.26 F1, 67.87% accuracy).
5.5. Model Performance Analysis
Figure 9 presents a comprehensive comparison of three distinct methodological approaches across six critical performance dimensions for multilingual receipt understanding. The Figure 9 reveals complementary strengths and trade-offs between traditional methods, advanced neural networks, and large language models. Traditional approaches (Tesseract + CAM) excel in real-time performance and show reasonable multilingual support but demonstrate significant limitations in object detection, OCR accuracy, and information extraction capabilities. Advanced neural networks (DINO + Azura) achieve superior performance in object detection and OCR accuracy with strong overall robustness, though at the cost of reduced real-time processing speed. LLM-based approaches demonstrate the highest information extraction capabilities and excellent multilingual support but suffer from poor real-time performance due to computational requirements.
Figure 9. Comprehensive performance comparison of traditional, advanced neural, and LLM-based approaches across six key dimensions for multilingual receipt understanding.
6. Conclusion
In this work, we presented ReceiptSense, a comprehensive dataset designed to advance multilingual OCR and post-OCR parsing, particularly for Arabic and English receipts. The dataset consists of over 20,000 annotated receipts, 30,000 OCR-annotated images, and 10,000 item-level annotations, supporting tasks such as object detection, OCR, and information extraction. We evaluated various models, ranging from traditional methods like Tesseract OCR to advanced neural network architectures like YOLO and DINO, as well as large language models like LLaMA and Mistral. Our results demonstrate the dataset’s effectiveness in capturing real-world receipt variations and highlight the challenges posed by noisy, multilingual text. By publicly releasing ReceiptSense, we aim to foster research into document understanding across diverse domains and applications.
7. GenAI Usage Disclosure
We used OpenAI’s ChatGPT for minor language editing, specifically to rephrase sentences and correct grammatical errors.